OpenStudy (anonymous):

If (-1, y) lies on the graph of y = 3^x+1, then y = 0 1/3 1

OpenStudy (isaiah.feynman):

@xo_kansasprincess_xo How to solve this, when x = -1, what is y?

OpenStudy (hockeychick23):

If (-1, y) lies on the graph of y = 3^x+1, then y = Ok so first we need to plug in -1 for x: y=3^-1+1. So 3^-1= 1/3+1. So 1/3+1=y y=1 1/3 Its not one of the answers you posted but its the correct answer

OpenStudy (anonymous):

@hockeychick23 That's obviously not the correct answer since it's not one of the choices!

OpenStudy (hockeychick23):

it's correct, whether its written as 4/3 or 1 1/3 and its hard to believe you have just 3 answer choices

OpenStudy (tkhunny):

@xo_kansasprincess_xo You HAVE to start showing your own work. This would have cleared up the confusion on this one. I don't believe you cannot substitute x = -1 into your equation. Your equation was mistaken for this \(y = 3^{x} + 1\). For this, the correct answer is y = 4/3, whether or not this is in the set of given answer choices. Had you written, y = 3^(x+1), it would have meant \(y = 3^{x+1}\), and there would have been no confusion. Now, go ahead and substitute x = -1 and resolve the right hand side to produce the corresponding y-value.